# Zero Trust Architecture: A Systematic Literature Review

### v1

@@ 6.1.4 Real-time Authentication and Authorization
One of the key requirement and feature of ZTA is continuous authentication of users, devices, and services [167], and dynamic authorization for network resources [166], [122]. However, meeting these requirements in real-time can present several issues and challenges including user experience and consumption of resources which can lead to performance degradation [103], especially for large and diverse networks. Additionally, the network is susceptible to increased latency [73] and reliability issues which some networks may not be able to tolerate depending on their real-time criticality. Real-time authentication and authorization is a post-implementation issue and proper approach must be considered to effectively harnessed the true power of ZTA.
@@ 6.4.2 Lightweight Cryptographic Solutions
While traditional cryptographic techniques such as AES, RSA, and Elliptic Curve Cryptographic (ECC) work on conventional computing devices like desktops, laptops, e.t.c, they are inadequate for IoT devices and embedded systems [6]. To accommodate the resource-constraint nature of IoT devices while ensuring their security, lightweight cryptographic algorithms are often proposed [189], [5]. However, these solutions may trade off robustness for efficiency, potentially introducing vulnerabilities. The challenge lies in balancing the need for minimal computational overhead with maintaining strong security guarantees required by ZTA [6].
@@ 6.5.2 Policy Coordination Challenges
Today’s enterprise networks are highly heterogeneous and dynamic especially with the proliferation of IoT and smart devices, and this has made management of security functions already complex due to huge array of policies [8]. Moreover, the development and management of ZT security rules can only get extremely complex and complicated in such networks since various applications and services would need different security policies [6], [45]. This is due to the fact that the security policies need to take into account each device’s authentication and authorization while safeguarding its data and communication across the system. Similarly, while features such as micro-segmentation can break large networks into smaller segments, the management of the individual segments with each having its own security requirements and policies can be a daunting task to execute [3], [4].
@@ 7 Conclusion
Through this SLR, we have provided a comprehensive analysis of ZTA, including its evolution and influencing factors, diverse application domains, enabling technologies, and the challenges associated with its adoption. The study contributes to the body of knowledge by offering a structured taxonomy and presents opportunities for future work to refine ZTA implementation strategies taking insights from existing frameworks proposed across several sectors, integrating advanced technologies, and addressing the identified barriers to adoption to maximize its potential in safeguarding critical infrastructures. While the trend of research on ZTA has seen an exponential surge in recent years with publications almost doubling that of the preceding year, the rate of growth has slowed down between 2023 and 2024 which may indicates that research interest on its applications maybe reaching a plateau. Also, with many studies proposing ZTA-based security frameworks especially in securing IoT networks and devices, there is lack of research on the core aspects of ZTA itself, and its associated weaknesses and trade-offs. Questions about ZTA’s perceived real-world security benefits versus its complexity, resource overhead and scalability, and implications on user experience remain largely unanswered. Similarly, the issue of trust management which is a critical aspect of ZTA, privacy preservation, and compliance with regulations need extensive exploration. Overall, by serving as a reference point for researchers and practitioners, this review aims to advance the effective implementation of ZTA and its role in fortifying cybersecurity in an increasingly interconnected world.

### v2

@@ 6.1.4 Real-time Authentication and Authorization
One of the key requirement and feature of ZTA is continuous authentication of users, devices, and services [173], and dynamic authorization for network resources [172], [128]. However, meeting these requirements in realtime can present several issues and challenges including user experience and consumption of resources which can lead to performance degradation [109], especially for large and diverse networks. Additionally, the network is susceptible to increased latency [79] and reliability issues which some networks may not be able to tolerate depending on their real-time criticality. Real-time authentication and authorization is a post-implementation issue and proper approach must be considered to effectively harnessed the true power of ZTA.
@@ 6.4.2 Lightweight Cryptographic Solutions
While traditional cryptographic techniques such as AES, RSA, and Elliptic Curve Cryptographic (ECC) work on conventional computing devices like desktops, laptops, e.t.c, they are inadequate for IoT devices and embedded systems [6]. To accommodate the resource-constrained nature of IoT devices while ensuring their security, lightweight cryptographic algorithms are often proposed [195], [5]. However, these solutions may trade off robustness for efficiency, potentially introducing vulnerabilities. The challenge lies in balancing the need for minimal computational overhead with maintaining strong security guarantees required by ZTA [6].
@@ 6.5.2 Policy Coordination Challenges
Today’s enterprise networks are highly heterogeneous and dynamic especially with the proliferation of IoT and smart devices, and this has made management of security functions already complex due to huge array of policies [8]. Moreover, the development and management of ZT security rules can only get extremely complex and complicated in such networks since various applications and services would need different security policies [6], [45]. This is due to the fact that the security policies need to take into account each device’s authentication and authorization while safeguarding its data and communication across the system. Similarly, while features such as micro-segmentation can break large networks into smaller segments, the management of the individual segments with each having its own security requirements and policies can be a daunting task to execute [3], [4]
@@ 7 Conclusion
Through this SLR, we have provided a comprehensive analysis of ZTA, including its evolution and influencing factors, diverse application domains, enabling technologies, and the challenges associated with its adoption. The study contributes to the body of knowledge by offering a structured taxonomy and presents opportunities for future work to refine ZTA implementation strategies taking insights from existing frameworks proposed across several sectors, integrating advanced technologies, and addressing the identified barriers to adoption to maximize its potential in safeguarding critical infrastructures. While the trend of research on ZTA has seen an exponential surge in recent years with publications almost doubling that of the preceding year, the rate of growth has slowed down between 2023 and 2024 which may indicates that research interest on its applications maybe reaching a plateau. Also, with many studies proposing ZTA-based security frameworks especially in securing IoT networks and devices, there is lack of research on the core aspects of ZTA itself, and its associated weaknesses and trade-offs. Questions about ZTA’s perceived real-world security benefits versus its complexity, resource overhead and scalability, and implications on user experience remain largely unanswered. Similarly, the issue of trust management which is a critical aspect of ZTA, privacy preservation, and compliance with regulations need extensive exploration. Moreover, study on various AI/ML techniques and other enabling technologies such as SASE, identity management techniques, etc, that will effectively help in realizing the requirements of ZTA can be further explored. Overall, by serving as a reference point for researchers and practitioners, this review aims to advance the effective implementation of ZTA and its role in fortifying cybersecurity in an increasingly interconnected world.